12. In one hour a car can make 600 buttons or 720 buttons. At 3 o'clock in the afternoon the car starts working. Makes 900 buttons and then changes to button production. How many buttons will the machine make by 8 p.m.?


1
Expert's answer
2021-05-31T19:12:02-0400

A car makes 900 buttons A in 900/600 = 1.5 hours.

The time left for making buttons B is:

8 - 3 - 1.5 = 3.5 hours.

So, the machine will produce 720×3.5 = 2,520 buttons B by 8 p.m.